# UrbanGraph: Physics-Informed Spatio-Temporal Dynamic Heterogeneous Graphs for Urban Microclimate Prediction

### v1

@@ Abstract
With rapid urbanization, predicting urban microclimates has become critical, as it affects building energy demand and public health risks. However, existing generative and homogeneous graph approaches fall short in capturing physical consistency, spatial dependencies, and temporal variability. To address this, we introduce UrbanGraph, a physics-informed framework integrating heterogeneous and dynamic spatio-temporal graphs. It encodes key physical processes—vegetation evapotranspiration, shading, and convective diffusion—while modeling complex spatial dependencies among diverse urban entities and their temporal evolution. We evaluate UrbanGraph on UMC4/12, a physics-based simulation dataset covering diverse urban configurations and climates. Results show that UrbanGraph improves R² by up to 10.8% and reduces FLOPs by 17.0% over all baselines, with heterogeneous and dynamic graphs contributing 3.5% and 7.1% gains. Our dataset provides the first high-resolution benchmark for spatio-temporal microclimate modeling, and our method extends to broader urban heterogeneous dynamic computing tasks.
footnotetext: Preprint. Under review at ICLR 2026.
@@ 1 Introduction
Although data-driven methods are promising, they still face challenges in accurately modeling the underlying physical processes. Station-based methods can only predict time series at discrete locations, neglecting spatial relationships and failing to generate continuous physical fields. Grid-based generative models are constrained by local receptive fields, which makes it difficult for them to capture long-range spatial dependencies (Carter et al., 2016; Kemppinen et al., 2024). Graph Neural Networks (GNNs) offer a more natural framework for modeling the spatial dependencies among urban entities. However, existing GNN-based approaches often lack physical consistency. They typically employ a uniform message-passing mechanism that cannot distinguish between different physical processes, such as vegetation evapotranspiration and building shading (Zhao et al., 2021). Furthermore, these methods struggle to model temporal variability. They mostly rely on a fixed graph structure, which is incapable of representing how physical processes evolve in real-time in response to changing environmental conditions. Consequently, there is a pressing need in the field for a unified framework capable of explicitly modeling multiple physical processes and their temporal evolution.
@@
Addressing this gap requires solving the following two core and orthogonal technical challenges. (i) A method must be designed to encode multiple, independent, and time-varying physical processes into a dynamic, heterogeneous graph structure. The fundamental difficulty lies in abstracting continuous physical fields (such as radiative transfer and fluid dynamics) into a discrete and computationally efficient graph topology, without losing critical physical information. (ii) A neural network architecture must be designed that can effectively process this complex sequence of graphs. This is highly challenging because the model must not only handle the dual dynamics of both node features and graph topology but also remain sensitive to the diverse types of relationships, each driven by different physical processes. Ultimately, the entire framework must strike a balance between physical interpretability and computational efficiency, while avoiding oversimplification of the urban morphology (Heo et al., ).
To address these challenges, we propose the UrbanGraph framework,, illustrated in Figure 1. It first employs a physics-informed graph representation method to explicitly map multiple physical processes, such as shading and convection, into different types of dynamic edges, thereby transforming physical laws into a computable graph structure. Subsequently, we design a spatio-temporal graph network centered on a heterogeneous message-passing mechanism. This mechanism can assign dedicated learning functions to different physical processes (Schlichtkrull et al., 2017) and capture their dynamic evolution. The effectiveness of this approach is validated on a custom high-fidelity, physics-based simulation benchmark dataset. The results on prediction accuracy and computational efficiency across multiple targets suggest that this framework can be generalized to other urban computing tasks.
Figure 1: The UrbanGraph framework. (a) Overall pipeline: Geospatial data is converted into a graph structure, which is processed by a spatio-temporal GNN with climate and time characteristics to generate high-resolution predictions. (b) The physics-informed dynamic graph concept: Nodes represent urban entities, while edges—representing physical interactions like building shading and convective diffusion—are dynamically reconfigured over time to reflect changing environmental conditions.
@@
  • We propose a physics-informed graph representation that explicitly encodes time-varying physical processes and static spatial relationships into the topology of a dynamic heterogeneous graph. This method offers a novel pathway for injecting temporally evolving domain knowledge into graph learning.
  • We develop a dynamic heterogeneous graph neural network architecture for the UrbanGraph framework, which efficiently learns from the complex graph sequences we propose through a heterogeneous message-passing mechanism. Comprehensive experiments demonstrate that the architecture achieves state-of-the-art performance in both prediction accuracy and computational efficiency. Compared to four categories of baselines, it improves accuracy by up to 10.8% (in R²) and efficiency by 17.0% (in FLOPs).
@@ 2 Related Work
Classical Microclimate Prediction Methods. Classical methods for urban microclimate prediction can be broadly categorized into two types. The first consists of physics-based simulation models, such as CFD and ENVI-met (Toparlar et al., 2017; Tsoka et al., 2018; Liu et al., 2021; Barros Moreira de Carvalho & Bueno da Silva, 2024). These methods offer high physical fidelity but suffer from immense computational overhead, making them impractical for large-scale, long-term time-series prediction tasks. The second category comprises data-driven approaches, including traditional machine learning (Arulmozhi et al., 2021; Alaoui et al., 2023) and grid-based deep learning models like CNNs (Kumar et al., 2021; Kastner & Dogan, 2023; Fujiwara et al., 2024). While these methods are computationally efficient, the former struggle to capture complex spatial dependencies, and the latter are constrained by the Euclidean data assumption, making them unable to process the inherently non-structural geometry of urban environments. To address these limitations, GNNs provide a more suitable framework, as their message-passing mechanism can directly model complex spatial dependencies (Kipf & Welling, 2016; Xu et al., 2019; Zhou et al., 2020).
Physics-Informed Methods. The physics-informed approach aims to enhance the learning process by incorporating fundamental knowledge and governing physical laws (Karniadakis et al., 2021). In the field of urban computing, this is often achieved either by adding the residuals of physical equations to the loss function as a soft constraint, known as a learning bias (Shao et al., 2023; Taghizadeh et al., 2025), or by designing specific network modules or message-passing mechanisms to simulate physical processes, known as an inductive bias (Xue, ; Qu et al., 2023; Gao et al., 2024). However, the former approach often increases training overhead due to the need to compute residuals of partial differential equations (PDEs), while the latter, despite imposing hard constraints, can sacrifice model flexibility. For tasks where the physical processes and rules are relatively well-defined, a more efficient and decoupled pathway is to modify the input data, an approach known as an observational bias (Banerjee et al., 2025). For example, Pan et al. (2025) embed principles of traffic flow physics into a step-wise framework for intersection flow prediction by extracting physical performance indicators as input features.
@@
Heterogeneous Graph Methods. In the context of urban physical field prediction, Graph Neural Networks typically simplify the complex urban system into a homogeneous graph (Yu et al., 2024; Zheng & Lu, 2024). However, this simplification limits the model’s fidelity and interpretability. Heterogeneous graphs, which consist of multiple types of nodes and edges, can represent the rich semantic relationships in complex systems (Schlichtkrull et al., 2017; Zhang et al., 2019; Zhao et al., 2021). By designing type-aware message-passing mechanisms, Heterogeneous Graph Neural Networks (HGNNs) have achieved success in various tasks, such as quantifying road network homogeneity (Xue et al., 2022), perceiving urban spatial heterogeneity (Xiao et al., 2023), learning urban region representations (Kim & Yoon, 2025), predicting the interactive behaviors of traffic participants (Li et al., 2021), and uncovering the dynamics of building carbon emissions (Yap et al., 2025). Although HGNNs have shown great potential in the field of urban computing, their application in microclimate prediction—specifically, leveraging them to explicitly differentiate between multiple physical processes to enhance model fidelity—remains an unexplored area.
@@
Dynamic Graph Methods. In applications for urban physical field prediction, Graph Neural Networks often rely on a static graph topology to represent the spatial relationships between entities (Mandal & Thakur, 2023; Shao et al., 2024; Xu et al., 2024). However, this assumption conflicts with physical reality, as the scope and intensity of physical processes (e.g., building shading) are determined in real-time by external environmental factors (e.g., solar position). Dynamic Graph Neural Networks (DGNNs) provide a more realistic framework for this problem (Skarding et al., 2020; Zheng et al., 2024). DGNNs have become a mainstream and effective approach for handling other urban tasks with time-varying interactions, particularly in traffic forecasting, demonstrating their potential in the field of urban computing (Zhao et al., 2020; Xie et al., 2020; Bui et al., 2022). In these mainstream applications, the evolution of the graph is typically treated as a data-driven, observational phenomenon (Li et al., 2019; Jin et al., 2020). In physical field prediction tasks, however, the graph topology (e.g., shading relationships) is explicitly reconfigured at each timestep by exogenous physical first principles. This fundamental difference gives rise to the need for a new class of DGNNs capable of learning from graph topologies that are actively reconfigured by physical first principles.
@@ 3 Preliminary
Problem Formulation. We model the urban environment by discretizing Geographic Information System (GIS) data into grid cells, where each cell is treated as a node $v\in{\mathbb{V}}$. The state of the environment is represented by a sequence of dynamic heterogeneous graphs $\{{\mathcal{G}}_{t}\}$, where the graph at timestep t is defined as ${\mathcal{G}}_{t}=({\mathbb{V}},\mathcal{E}_{t},{\mathbb{R}})$. Here, $\mathbb{V}$} is the static set of nodes, $\mathbb{R}$} is the static set of relation types (e.g., ’covered by shadow from cell’), and $\mathcal{E}_{t}$ is the set of edges that varies with time. The features of all nodes are collected in a matrix ${\bm{X}}\in\mathbb{R}^{|{\mathbb{V}}|\times 8}$. Additionally, we define ${\bm{u}}_{t}$ as the global context vector at timestep t.
@@ 4 Method
Our proposed UrbanGraph framework consists of two core components: a physics-informed graph representation and a spatio-temporal dynamic relational graph network. To rigorously evaluate the effectiveness of our approach, we first generated a large-scale spatio-temporal dataset through high-fidelity physical simulations, the detailed generation process and parameter configurations of which are described in Appendix B. Second, to address the challenge that urban systems exhibit high heterogeneity in both spatial and temporal dimensions, we detail our physics-informed graph representation in Section 4.1, which is designed to efficiently capture the underlying physical interactions among different urban elements. Finally, in Section 4.2, we introduce the UrbanGraph architecture, which explicitly leverages the time-varying relationships between urban elements to perform node prediction tasks.
@@
Relational graph convolutional networks. This category of edges represents spatial and semantic relationships that do not change over time. As demonstrated by Yan et al. (2021), non-local spatial functions can significantly influence the assessment of a central region. To capture these non-local interactions between functionally similar nodes, we introduce the following edge type:
Semantic Similarity Edges.
Using the k-Nearest Neighbors (k-NN) algorithm, we construct directed edges from each node to its k nearest neighbors in the normalized static feature space.
Internal Contiguity Edges.
To simulate local energy transfer within large continuous bodies (e.g., building clusters or groups of trees), ’internal nodes’ establish connections with their eight immediate neighbors (Moore neighborhood). A node is defined as an ’internal node’ if and only if all of its direct neighbors (von Neumann neighborhood) belong to the same object class as itself.
@@
This edge type is used to model the cooling effect of shadows. A directed edge of type ’shadow’ is established from a shading object node $v_{i}$ (building or tree) with height $h_{obj}$ to a ground node $v_{j}$ if their Euclidean distance $d(v_{i},v_{j})$ is less than or equal to the shadow length $L_{shadow,t}$, and the angular deviation falls within a predefined shadow angle width $\Delta\varphi_{shadow}$. The shadow properties are calculated as follows:
@@
This edge type is designed to represent the local cooling effect of vegetation. A directed edge of type ’Vegetation Evapotranspiration’ is established from a tree node $v_{i}$ to any other node $v_{j}$ if their Euclidean distance $d(v_{i},v_{j})$ does not exceed a dynamic radius of influence, $R_{activity,t}$. This radius is calculated based on the global horizontal radiation $I_{t}$(Wh/m2) for the current hour, where $R_{base}$ is a presettable base radius:
@@ 4.2 UrbanGraph Architecture
To effectively process the graph structures from our physics-informed representation, we designed a dynamic and heterogeneous architecture for UrbanGraph. As illustrated in Figure 3, the overall architecture comprises four core components: Feature Encoders, a Spatial Graph Encoder, a Spatio-Temporal Evolution Module, and a Prediction Head.
Figure 3: The overall architecture of the UrbanGraph model.(a) The end-to-end framework, which processes historical time-series data, weather context, and a sequence of dynamic heterogeneous graphs. At each timestep, an RGCN Block extracts spatial features from the corresponding graph, which are then fused with temporal features by an MLP Fusion Layer. An LSTM layer propagates the temporal state to the next timestep.(b) The detailed structure of the RGCN Block, which aggregates messages from neighbors across different relation types (R1R5R1-R5) and combines them with the node’s self-features.(c) The architecture of a standard LSTM layer used for capturing temporal dependencies.
Feature Encoders and Spatial Graph Encoder. At each timestep t, we employ independent Multi-Layer Perceptrons (MLPs) to encode non-graph dynamic inputs. Graph-level global environmental features ${\bm{u}}_{t}^{env}$ and periodic temporal features ${\bm{u}}_{t}^{time}$ are projected into high-dimensional embedding vectors ${\bm{e}}_{t}^{env}$ and ${\bm{e}}_{t}^{time}$, respectively. For each graph ${\mathcal{G}}_{t}$ in the input sequence, we utilize a three-layer RGCN to capture the spatial dependencies defined by the time-varying heterogeneous edges. This module outputs a spatially-informed representation vector ${\bm{h}}_{v,t}^{RGCN}$ for each node $v$.
Spatio-Temporal Evolution Module. This module is responsible for fusing the spatial and global dynamic features and uses a Long Short-Term Memory (LSTM) network to model their temporal evolution. At each prediction timestep t (from $t_{1}$ to $T_{pred}$), we concatenate the spatial representation of a node, ${\bm{h}}_{v,t}^{RGCN}$, with the global environmental embedding, ${\bm{e}}_{t}^{env}$, and the temporal embedding, ${\bm{e}}_{t}^{time}$. The resulting concatenated vector is passed through a fusion MLP to generate the input feature for the LSTM layer, ${\bm{x}}_{v,t}^{LSTM}$. This is expressed as:

### v2

@@ Abstract
With rapid urbanization, predicting urban microclimates has become critical, as it affects building energy demand and public health risks. However, existing generative and homogeneous graph approaches fall short in capturing physical consistency, spatial dependencies, and temporal variability. To address this, we introduce UrbanGraph, a framework founded on a novel structure-based inductive bias. Unlike implicit graph learning, UrbanGraph transforms physical first principles into a dynamic causal topology, explicitly encoding time-varying causalities (e.g., shading and convection) directly into the graph structure to ensure physical consistency and data efficiency. Results show that UrbanGraph achieves state-of-the-art performance across all baselines. Specifically, the use of explicit causal pruning significantly reduces the model’s floating-point operations (FLOPs) by 73.8% and increases training speed by 21% compared to implicit graphs. Our contribution includes the first high-resolution benchmark for spatio-temporal microclimate modeling, and a generalizable explicit topological encoding paradigm applicable to urban spatio-temporal dynamics governed by known physical equations.
@@ 1 Introduction
Although data-driven methods are promising, they still face challenges in accurately modeling the underlying physical processes. While urban data is spatially discretized, physical interactions are often non-local (e.g., shadows skipping intermediate spaces) and anisotropic (e.g., directional wind flow). Grid-based models (CNNs) struggle to capture these long-range, irregular dependencies without excessive depth.(Carter et al., 2016; Kemppinen et al., 2024). Graph Neural Networks (GNNs) offer a more natural framework for modeling the spatial dependencies among urban entities. However, existing GNN-based approaches often lack physical consistency. They typically employ a uniform message-passing mechanism that cannot distinguish between different physical processes, such as vegetation evapotranspiration and building shading (Zhao et al., 2021). Furthermore, these methods struggle to model temporal variability. They mostly rely on a fixed graph structure, which is incapable of representing how physical processes evolve in real-time in response to changing environmental conditions. Consequently, there is a pressing need in the field for a unified framework capable of explicitly modeling multiple physical processes and their temporal evolution.
@@
Addressing this gap necessitates a fundamental shift in how we model urban dynamics. The core challenge lies in designing a structure-based inductive bias capable of encoding multiple, independent, and time-varying physical processes. (i) Standard graph topologies struggle to abstract continuous physical fields (e.g., radiative transfer and fluid dynamics) into a discrete representation without losing critical causal information. (ii) Furthermore, processing such complex graph sequences requires a neural architecture that can disentangle the diverse physical interactions. The model must not only handle the dual dynamics of both node features and graph topology but also differentiate between distinct physical operators (e.g., shading vs. convection), striking a balance between physical interpretability and computational efficiency(Heo et al., ).
To address these challenges, we propose UrbanGraph (see Figure 1), a framework founded on a novel structure-based inductive bias for spatio-temporal modeling. Unlike standard approaches that rely on implicit latent graph learning, UrbanGraph transforms physical first principles into a dynamic causal topology. By explicitly encoding time-varying causalities—such as solar shading and anisotropic wind convection—directly into the graph structure, we impose a hard structural constraint that forces the model’s receptive field to align with the actual physical domain of influence. This explicit causal encoding effectively reduces the hypothesis space, preventing the model from learning spurious correlations from noise. Subsequently, we design a spatio-temporal graph network where the heterogeneous message-passing mechanism functions as a physical operator approximator, assigning dedicated learnable parameters to disentangle distinct physical processes(Schlichtkrull et al., 2017).
Figure 1: Overview of the UrbanGraph framework. (a) Spatio-Temporal Pipeline: The model transforms rasterized urban features and dynamic weather conditions into high-resolution microclimate heatmaps via GNN learning. (b) Physics-Informed Topology Construction: Physical laws (e.g., shading, diffusion) serve as hard structural constraints to dynamically prune non-causal edges from the candidate set, constructing a sparse and physically consistent graph topology.
@@
  • We propose a structure-based inductive bias via dynamic topological reconfiguration. By explicitly encoding time-varying physical processes into the topology of a dynamic heterogeneous graph, this method offers a novel pathway for injecting time-evolving causal knowledge into graph learning as a hard constraint.
  • We develop a dynamic heterogeneous graph neural network architecture that efficiently learns from complex graph sequences. Through a specialized heterogeneous message-passing mechanism, the model achieves physical operator decoupling, allowing for precise modeling of distinct environmental interactions.Comprehensive experiments demonstrate that the architecture achieves state-of-the-art performance in both prediction accuracy and computational efficiency. Compared to four categories of baselines, it improves accuracy by up to 10.8% (in R²) and efficiency by 17.0% (in FLOPs).
@@ 2 Related Work
Classical Microclimate Prediction Methods. Classical methods for urban microclimate prediction can be broadly categorized into two types. The first consists of physics-based simulation models, such as CFD and ENVI-met (Toparlar et al., 2017; Tsoka et al., 2018; Liu et al., 2021; Barros Moreira de Carvalho and Bueno da Silva, 2024). These methods offer high physical fidelity but suffer from immense computational overhead, making them impractical for large-scale, long-term time-series prediction tasks. The second category comprises data-driven approaches, including traditional machine learning (Arulmozhi et al., 2021; Alaoui et al., 2023) and grid-based deep learning models like CNNs (Kumar et al., 2021; Kastner and Dogan, 2023; Fujiwara et al., 2024). While these methods are computationally efficient, the former struggle to capture complex spatial dependencies, and the latter are constrained by the Euclidean data assumption, making them unable to process the inherently non-structural geometry of urban environments.While generic GNNs (Kipf and Welling, 2016; Xu et al., 2019; Zhou et al., 2020) offer a framework for spatial dependencies, they lack the inductive bias to abstract continuous fields. UrbanGraph addresses this by explicitly bridging the gap between continuous field dynamics and discrete graph topology.
Physics-Informed Methods. The physics-informed approach incorporates fundamental knowledge into the learning process (Karniadakis et al., 2021). This integration is achieved primarily through three bias types. The learning bias uses soft constraints (e.g., PDE residuals in the loss function) but incurs significant training overhead (Shao et al., 2023; Taghizadeh et al., 2025). The observational bias leverages physics to modify input features (e.g., by extracting high-level physical indicators) (Pan et al., 2025). A third type is the inductive bias, which imposes hard constraints via network modules or graph structures to simulate physical processes (Xue, ; Qu et al., 2023; Gao et al., 2024). However, this traditional hard-constraint approach often sacrifices model flexibility. UrbanGraph addresses this challenge by adopting a novel, efficient form of the inductive bias. By imposing physics as a dynamic structural constraint via graph topology, UrbanGraph ensures consistency and interpretability without sacrificing necessary flexibility or incurring the PDE solver overhead associated with the learning bias.
@@
Heterogeneous Graph Methods. In the context of urban physical field prediction, Graph Neural Networks typically simplify the complex urban system into a homogeneous graph (Yu et al., 2024; Zheng and Lu, 2024). However, this simplification limits the model’s fidelity and interpretability. Heterogeneous graphs, which consist of multiple types of nodes and edges, can represent the rich semantic relationships in complex systems (Schlichtkrull et al., 2017; Zhang et al., 2019; Zhao et al., 2021). By designing type-aware message-passing mechanisms, Heterogeneous Graph Neural Networks (HGNNs) have achieved success in various tasks, such as quantifying road network homogeneity (Xue et al., 2022), perceiving urban spatial heterogeneity (Xiao et al., 2023), learning urban region representations (Kim and Yoon, 2025), predicting the interactive behaviors of traffic participants (Li et al., 2021), and uncovering the dynamics of building carbon emissions (Yap et al., 2025). We bridge this gap by leveraging heterogeneity for physical operator decoupling, assigning distinct learnable operators to fundamentally different environmental interactions.
@@
Dynamic Graph Methods. In applications for urban physical field prediction, Graph Neural Networks often rely on a static graph topology to represent the spatial relationships between entities (Mandal and Thakur, 2023; Shao et al., 2024; Xu et al., 2024). However, this assumption conflicts with physical reality, as the scope and intensity of physical processes (e.g., building shading) are determined in real-time by external environmental factors (e.g., solar position). Dynamic Graph Neural Networks (DGNNs) provide a more realistic framework for this problem (Skarding et al., 2020; Zheng et al., 2024). DGNNs have become a mainstream and effective approach for handling other urban tasks with time-varying interactions, particularly in traffic forecasting, demonstrating their potential in the field of urban computing (Zhao et al., 2020; Xie et al., 2020; Bui et al., 2022). In these mainstream applications, the evolution of the graph is typically treated as a data-driven, observational phenomenon (Li et al., 2019a; Jin et al., 2020). In physical field prediction tasks, however, the graph topology (e.g., shading relationships) is explicitly reconfigured at each timestep by exogenous physical first principles. Consequently, we propose a First-Principle-Driven approach that utilizes dynamic topological reconfiguration as a causal pruning mechanism to explicitly model time-varying interactions.
@@ 3 Preliminary
Problem Formulation. We model the urban environment by discretizing Geographic Information System (GIS) data into grid cells, where each cell is treated as a node $v\in{\mathbb{V}}$. The state of the environment is represented by a sequence of dynamic heterogeneous graphs $\{{\mathcal{G}}_{t}\}$, where the graph at timestep t is defined as ${\mathcal{G}}_{t}=({\mathbb{V}},\mathcal{E}_{t},{\mathbb{R}})$. Here, $\mathbb{V}$} is the static set of nodes, $\mathbb{R}$} is the static set of relation types (e.g., ’covered by shadow from cell’), and $\mathcal{E}_{t}$ is the set of edges that varies with time. The static node feature matrix $\bm{X}\in\mathbb{R}^{|\mathbb{V}|\times 8}$ explicitly encodes pixel-wise GIS properties: Building Height, Tree Height and Land Cover Type. Additionally, $\bm{u}_{t}$ represents the dynamic global context, comprising meteorological forcing data: Solar Radiation, Solar Position, Ambient Temperature, Humidity, Wind Speed, and Wind Direction.
@@ 4 Method
Our proposed UrbanGraph framework consists of two core components: a physics-informed graph representation and a spatio-temporal dynamic relational graph network. To rigorously evaluate the effectiveness of our approach, we first generated a large-scale spatio-temporal dataset through high-fidelity physical simulations, the detailed generation process and parameter configurations of which are described in Appendix B.1. Second, to address the challenge that urban systems exhibit high heterogeneity in both spatial and temporal dimensions, we detail our physics-informed graph representation in Section 4.1, which is designed to efficiently capture the underlying physical interactions among different urban elements. Finally, in Section 4.2, we introduce the UrbanGraph architecture, which explicitly leverages the time-varying relationships between urban elements to perform node prediction tasks.
@@
This edge type encodes directional radiative obstruction. By establishing links based on geometric line-of-sight, we strictly enforce the dependency of the shadowed region on the occluding object. A directed edge of type ’shadow’ is established from a shading object node $v_{i}$ (building or tree) with height $h_{obj}$ to a ground node $v_{j}$ if their Euclidean distance $d(v_{i},v_{j})$ is less than or equal to the shadow length $L_{shadow,t}$, and the angular deviation falls within a predefined shadow angle width $\Delta\varphi_{shadow}$. The shadow properties are calculated as follows:
@@
This edge type models localized bio-physical interactions, serving as a solar-dependent spatial filter. Specifically, a directed edge is established from a tree node $v_{i}$ to any other node $v_{j}$ if their Euclidean distance $d(v_{i},v_{j})$ does not exceed a dynamic radius of influence, $R_{activity,t}$. This radius is calculated based on the global horizontal radiation $I_{t}$(Wh/m2) for the current hour, where $R_{base}$ is a presettable base radius:
@@
Static Semantic Topology. To complement the dynamic physical edges and provide structural redundancy against heuristic imperfections, we introduce two static edge types that capture non-local dependencies and local continuity:
Semantic Similarity Edges.
To capture non-local functional interactions, we construct directed edges from each node to its $k$ nearest neighbors in the normalized static feature space. This mechanism allows the GNN to aggregate information from spatially distant but physically similar entities (e.g., similar materials), serving as a backup information pathway.
Internal Contiguity Edges.
To model intra-object energy transfer (thermal inertia) within large continuous bodies, ’internal nodes’ establish connections with their eight immediate neighbors (Moore neighborhood). This ensures local physical consistency within building clusters or vegetation patches.
@@ 4.2 UrbanGraph Architecture
To align neural computation with the encoded structural priors, we designed a dynamic and heterogeneous architecture for UrbanGraph. As illustrated in Figure 3, the overall architecture comprises four core components: Feature Encoders, a Spatial Graph Encoder, a Spatio-Temporal Evolution Module, and a Prediction Head.
Figure 3: Architecture of UrbanGraph. (a) End-to-End Framework: Processes historical data, weather, and dynamic graphs. The top-left grids denote global and cyclically encoded time features. At each step, RGCN blocks extract spatial features, fused via MLP before LSTM temporal propagation. (b) RGCN Block: Aggregates multi-relational neighbor messages (R1R5R1-R5) with node self-features. (c) LSTM Layer: Captures temporal dependencies.
Feature Encoders and Spatial Graph Encoder. At timestep $t$, MLP-encoded global environmental ($u_{t}^{env}$) and temporal ($u_{t}^{time}$) features are broadcast and concatenated with spatial node representations $h_{v,t}^{RGCN}$ extracted by a three-layer RGCN. This fusion enables the subsequent LSTM to contextualize global forcing within local topologies. Crucially, the RGCN functions as a physical operator approximator: by assigning dedicated weight matrices $W_{r}$ (Eq. 2) to specific relations (e.g., shading vs. convection), it disentangles complex dynamics into distinct physical sub-processes, structurally mitigating the over-smoothing issues typical of homogeneous GNNs.
Spatio-Temporal Evolution Module. This module is responsible for fusing the spatial and global dynamic features and uses a Long Short-Term Memory (LSTM) network to model their temporal evolution. We specifically select LSTM over Attention mechanisms to align with the Markovian nature of physical transport processes (e.g., heat diffusion), where the future state evolves continuously from the immediate past. Furthermore, unlike coupled spatio-temporal operators (e.g., LRGCN), our architecture decouples spatial interaction from temporal evolution. By resolving the dynamic topology explicitly within the RGCN module, we provide the LSTM with a physics-consistent state representation, thereby reducing optimization complexity.
At each prediction timestep t (from $t_{1}$ to $T_{pred}$), we concatenate the spatial representation of a node, ${\bm{h}}_{v,t}^{RGCN}$, with the global environmental embedding, ${\bm{e}}_{t}^{env}$, and the temporal embedding, ${\bm{e}}_{t}^{time}$. The resulting concatenated vector is passed through a fusion MLP to generate the input feature for the LSTM layer, ${\bm{x}}_{v,t}^{LSTM}$. This is expressed as: